In a secret code, each letter stands for a single digit- 0,1,2,3,4,5,6,7,8,9. Different letters stand for different digits. Using the code, a certain addition problem works out correctly as follows:
S E V E N
+ E I G H T
------------
T W E L V E Assume that neither S nor E equals the digit 0, since no numbers begin with a 0.
S=___, E=___, V=__, N=___, I=__, G=___, H=___, T=___, W=___, L=___

Answer by Edwin McCravy(20065)

There are 6 different solutions:

    SEVEN + EIGHT = TWELVE 
1.  36465 + 69781 = 106246
2.  38487 + 89561 = 128048
3.  58287 + 80641 = 138928
4.  63732 + 39841 = 103573
5.  69298 + 90431 = 159729
6.  85254 + 50671 = 135925

Edwin